### v1

# Learning Linearity in Audio Consistency Autoencoders via Implicit Regularization

Abstract
Audio autoencoders learn useful, compressed audio representations, but their non-linear latent spaces prevent intuitive algebraic manipulation such as mixing or scaling. We introduce a simple training methodology to induce linearity in a high-compression Consistency Autoencoder (CAE) by using data augmentation, thereby inducing homogeneity (equivariance to scalar gain) and additivity (the decoder preserves addition) without altering the model’s architecture or loss function. When trained with our method, the CAE exhibits linear behavior in both the encoder and decoder while preserving reconstruction fidelity. We test the practical utility of our learned space on music source composition and separation via simple latent arithmetic. This work presents a straightforward technique for constructing structured latent spaces, enabling more intuitive and efficient audio processing.
Index Terms—  audio, compression, diffusion, source separation
1 Introduction
Modern Autoencoders can achieve excellent reconstruction quality at high compression rates at the expense of complex, entangled latent spaces. For applications where input space manipulation is desirable from within the compressed space, recent research has proposed either task-specific adaptors for pre-trained models or redesigning the autoencoder to preserve key structural properties, such as equivariance to spatial transformations [1, 2, 3]). This work follows the latter approach.
For certain applications in audio, linearity (Figure 1) is a desirable property. A linear map fulfills two properties : (I) Homogeneity: scaling the input by a value scales the output by the same value; and (II) Additivity: the map preserves addition. As processing large audio datasets in the latent space becomes more prevalent, direct mixing and volume adjustment in this space can improve efficiency by reducing redundant encoding and decoding. Additionally, as downstream tasks such as audio generation and source separation can benefit from composition via latent arithmetic, better interpretability may be achieved when working on a linear space.
This work presents a training methodology for constructing an approximately linear compressed audio representation that enables intuitive manipulation, where simple algebraic operations in the latent space correspond directly to mixing and scaling in the audio domain. The method employs implicit regularization through data augmentation, without modifying the model architecture or objective. The approach is demonstrated with the Music2Latent architecture [4], a Consistency Autoencoder (CAE) that achieves high-quality, single-step reconstruction and a $64\times$ compression rate for $44.1$ kHz audio.
Figure 1: In a linear decoder, applying a gain to the latent vector scales the output by the same gain (homogeneity), and summing latents corresponds to a sum in the audio domain (additivity).
Our main contributions are: (I) An unsupervised, data-augmentation-based training procedure that induces approximate linearity in a high compression AE, with no extra loss terms; (II) Validation on a state-of-the-art CAE for music and speech, showing linearity in both encoder and decoder with no loss of reconstruction quality; (III) Practical utility shown on oracle source separation via simple latent arithmetic. Code and model weights are available online111www.github.com/bernardo-torres/linear-autoencoders..
(a) Music2Latent
(b) CAE training with implicit homogeneity
(c) CAE training with implicit additivity
(d) Batch creation
Figure 2: (2(a)): Music2Latent CAE architecture. The decoder is a denoising U-Net and the latent is introduced to it at every resolution level after learned upsampling. (2(b)): Proposed CAE training trick to implicitly enforce homogeneity in the decoder. (2(c)): Proposed trick to enforce additivity, applied when the input is an artificial mixture. (2(d)): Batch creation procedure with artificial mixtures of mixtures.
2 Background
2.1 Audio processing in the latent space
There is a growing interest in manipulating audio by training task-specific modules which operate directly in the latent space of a pretrained AE, with applications in source separation [5], speech enhancement [6, 7], upsampling and upmixing [8], filtering [9], and generative modelling [10, 11, 12]. Some works retrain AEs for specific objectives, such as source disentanglement [13] or separation via latent masking [14]. In contrast, our work focuses on training an AE to have desirable, task-agnostic structural properties. This enables direct and efficient audio manipulation and can serve as a strong foundation for these downstream applications.
2.2 Diffusion and Consistency models
Denoising Diffusion Probabilistic Models [15] and score-based models [16] have achieved great success in generative modeling, where the goal is to estimate the underlying data distribution from samples. These models define a forward process that gradually adds noise to data and learn a reverse process to denoise it using a neural network. Sampling typically requires an iterative procedure with many steps to generate a clean sample. Consistency Models [17] accelerate this process by mapping any point along the trajectory defined by the probability flow ordinary differential equation [16] directly to the origin (the clean data point), enabling single-step generation. When trained from scratch, the process is called Consistency Training (CT), in which a ”student” denoiser network ($f_{\theta}$) is trained to match the output of a ”teacher” ($f_{\theta^{-}}$), which itself denoises a less corrupted version of the same data point. In improved Consistency Training (iCT) [18], the teacher is updated with the same parameters $\theta$ as the student but detached from the computational graph ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$).
2.3 High-Fidelity Audio Autoencoders
In audio, popular AEs include Neural Audio Codecs [19], which compress audio into a set of discrete tokens, and Variational Autoencoders [20], which have seen significant success in generative modeling. However, both frameworks require complex, multi-stage training with adversarial objectives [21] to achieve high-quality reconstructions. Diffusion Autoencoders [22, 23, 24, 25, 10] enable high-fidelity reconstruction by replacing the deterministic decoder with a conditional diffusion model. Notable examples in audio include [10] and Music2Latent [4, 26]. A specific instance of DiffAEs is the CAE[4, 26], where the decoder is a CM, enabling decoding in one step. DiffAEs can be trained with a single diffusion/CT objective [22, 24, 4]. High-quality decoding has been linked to the capacity of sampling details at inference-time, since the decoder acts as a denoiser instead of an upsampler, which also reduces the amount of unnecessary information to be encoded in the latent [23, 25].
3 Method
Let $\mathcal{M}\subset[-1,1]^{T}$ be the space of audio signals of length $T$ of interest and $\mathcal{Z}\subset{\mathbb{R}}^{N\times F}$ be a lower-dimensional latent space induced by encoder $\operatorname{Enc}_{\theta}:\mathcal{M}\to\mathcal{Z}$ and decoder $\operatorname{Dec}_{\theta}:\mathcal{Z}\to\mathcal{M}$, with dimensions $N,F$ defined by the compression factor. We train $(\operatorname{Enc}_{\theta},\operatorname{Dec}_{\theta}$) with CAE training under the constraint that $\operatorname{Dec}_{\theta}$ is approximately linear (Figure 1), i.e., it satisfies the following properties:
Property 1 (Homogeneity).
$\operatorname{Dec}_{\theta}({{\color[rgb]{1,0,1}a}}\cdot{\mathbf{z}}_{x})\approx{{\color[rgb]{1,0,1}a}}\cdot\operatorname{Dec}_{\theta}({\mathbf{z}}_{x}),\quad\forall{\mathbf{z}}_{x}\in\mathcal{Z},{{\color[rgb]{1,0,1}a}}\in{\mathbb{R}}$ (1)
Property 2 (Additivity).
$\operatorname{Dec}_{\theta}({\mathbf{z}}_{u}+{\mathbf{z}}_{v})\approx\operatorname{Dec}_{\theta}({\mathbf{z}}_{u})+\operatorname{Dec}_{\theta}({\mathbf{z}}_{v}),\quad\forall{\mathbf{z}}_{u},{\mathbf{z}}_{v}\in\mathcal{Z}$ (2)
We postulate that approximate linearity can be achieved by a simple data augmentation scheme, without the need to change model or loss function. Our approach is in theory model-agnostic and can be applied to any AE, and in this work we apply it on a CAE architecture for audio [4], which offers $64\times$ compression and high-quality reconstruction with a single training objective.
3.1 Model architecture
Our model (Lin-CAE) follows the Music2Latent architecture [4], illustrated in Figure 2(2(a)). We recall the main components here for clarity, but we refer the reader to [4] for the full details.
CAE Denoiser representation space: We use a complex-valued Short-Time Fourier Transform (STFT) followed by the invertible amplitude scaling ${\operatorname{Amp}}:{\mathbb{C}}\to{\mathbb{C}}=\beta|c|^{\alpha}e^{i\angle(c)}$ (with $\alpha=0.65$ and $\beta=0.34$) from previous work in complex STFT diffusion [27, 28, 4], which scales the amplitudes in the STFT to roughly [1,1][-1,1] while boosting the high frequencies. We define ${\operatorname{T}}(x)={\operatorname{Amp}}(\operatorname{STFT}(x))$ as the transform which maps from waveform to CM input space.
Decoder: $\operatorname{Dec}_{\theta}$ is composed of a U-Net-based denoiser $f_{\theta}$} and the inverse transform ${\operatorname{T}}^{-1}({\mathbf{x}})$. The denoiser reconstructs ${\mathbf{x}}={\operatorname{T}}(x)$ from a noise-corrupted version ${\mathbf{x}}_{\sigma}={\mathbf{x}}+\sigma\bm{\epsilon}$ ($\bm{\epsilon}\sim\mathcal{N}(0,I)$), conditioned on the latent ${\mathbf{z}}_{x}$ and the noise level $\sigma$. ${\mathbf{z}}_{x}$ is upsampled with a dedicated network mirroring the U-Net’s upsampling block and added to each layer of the U-Net. Noise level information is added to every layer via positional embeddings. $f_{\theta}$} is parameterized by a noise prediction network $F_{\theta}$} with a skip connection weighted by coefficients ($c_{\text{skip}},c_{\text{out}}$) that depend on $\sigma$ [29], which additionally enforces the boundary condition necessary for CT [17, 18, 4]:
${f_{\theta}}({\mathbf{x}}_{\sigma},\sigma,{\mathbf{z}}_{x})=c_{\text{skip}}(\sigma){\mathbf{x}}_{\sigma}+c_{\text{out}}(\sigma){F_{\theta}}({\mathbf{x}}_{\sigma},\sigma,{\mathbf{z}}_{x})$ (3)
During inference, decoding starts from pure noise and the clean signal is reconstructed in a single step conditioned on ${\mathbf{z}}_{x}$.
Encoder: $\operatorname{Enc}_{\theta}$ consists of the amplitude transform $\operatorname{T}$} followed by a network that mirrors the U-Net’s downsampling blocks.
3.2 Learning Linearity in CAEs via Implicit Regularization
Implicit homogeneity: Inspired by [1], we apply a random positive gain ($\color[rgb]{1,0,1}a$}}) to the latent and task the decoder to reconstruct a scaled version of the input (${{\color[rgb]{1,0,1}a}}x$). The adapted CT training, depicted in Figure 2(2(b)), uses two parallel pathways. $\operatorname{Enc}_{\theta}$ encodes unscaled $x$ to obtain ${\mathbf{z}}_{x}$. Then, the denoiser ($f_{\theta}$}) receives a noisy scaled input (${\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma\bm{\epsilon}$) conditioned on the scaled latent (${{\color[rgb]{1,0,1}a}}{\mathbf{z}}_{x}$). $\color[rgb]{1,0,1}a$}} is never provided as an explicit input to any part of the model, so $\operatorname{Dec}_{\theta}$ must learn to infer the correct output scale from the magnitude of the conditioning latent. If ${{\color[rgb]{1,0,1}a}}=1$, we recover the original CAE training.
Implicit additivity: Inspired by Mixit [30], we augment our data by creating artificial mixes from pairs of elements $(u,v)$ randomly selected from the training set. When decoding, we replace the true latent of the artificial mix ($\operatorname{Enc}_{\theta}(u+v)$) with the sum of the latents of the individual signals: ${\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}}\leftarrow({\color[rgb]{0,0,1}{\mathbf{z}}_{u}+{\mathbf{z}}_{v}})$, where $({\mathbf{z}}_{u}=\operatorname{Enc}_{\theta}(u),{\mathbf{z}}_{v}=\operatorname{Enc}_{\theta}(v))$, as illustrated in Figure 2(2(c)). The denoiser is thus conditioned on $\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}$} and tasked with denoising ${\operatorname{T}}({{\color[rgb]{1,0,1}a}}(u+v))+\sigma\bm{\epsilon})$. We create mixtures from each training batch by concatenating it with a version of itself that has been circularly shifted by one position and summed with the original (Figure 2(2(d))).
3.3 Training
With the modifications described in Section 3.2 highlighted in color, we train the model under the CT objective $\mathcal{L}_{\text{CT}}$ [17, 18, 4], which minimizes the distance between the denoised output of a teacher ($f_{\theta^{-}}$) and student ($f_{\theta}$) models ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$) for noisy inputs at two different noise levels, $\sigma_{t_{1}}$ and $\sigma_{t_{2}}$.
$\begin{split}\mathcal{L}_{\text{CT}}=\mathbb{E}_{{\mathbf{x}},\bm{\epsilon},t_{1},{{\color[rgb]{1,0,1}a}}}\bigg[\lambda\cdot d\Big(f_{\theta}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma_{t_{2}}\bm{\epsilon},\sigma_{t_{2}},{{\color[rgb]{1,0,1}a}}{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime}),\\ f_{\theta^{-}}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma_{t_{1}}\bm{\epsilon},\sigma_{t_{1}},{{\color[rgb]{1,0,1}a}}{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime})\Big)\bigg],\end{split}$ (4)
where $\color[rgb]{0,0,1}{\mathbf{z}}$$^{\prime}$ is either ${\mathbf{z}}_{x}$ or ${\mathbf{z}}_{u}+{\mathbf{z}}_{v}$ depending on whether the input waveform is sampled from the training set or is an artificial mixture ($x=u+v$), $\bm{\epsilon}\sim\mathcal{N}(0,I)$ is a shared noise direction, $\lambda(\sigma_{t_{1}},\sigma_{t_{2}})=\frac{1}{\sigma_{t_{2}}-\sigma_{t_{1}}}$ is a weighting function designed to give higher weights to smaller noise increments [18]. We sample time steps $t_{1}\in[0,1]$ and corresponding noise levels $\sigma_{t_{1},t_{2}}$ following [4], with $t_{2}=t_{1}+\Delta_{t_{k}}$, where $\Delta_{t_{k}}$ depends on the training step $k$ and decays over training ($\Delta t_{k}=\Delta t_{0}^{\frac{k}{K}(e_{K}-1)+1}$, $\Delta t_{0}=0.1$, $e_{K}=2.0$). $d$ is the Pseudo-Huber loss [31]: $d(x,y)=\sqrt{|x-y|^{2}+c^{2}}-c$, with c=5.4e4c=5.4e-4.
Training details: We train all models for 800800K steps with a batch size of $20$ before mixture creation (final batch size of $40$). Training takes $\approx$ 88 days on 11 L4040S GPU. The models are optimized with RAdam with learning rate $\in[10^{-4},10^{-6}]$ (linear warmup for 1010K steps, cosine decay for the rest). We track an Exponential Moving Average (EMA) of the model parameters $\theta$ to be used for inference, updated every $10$ steps.
Homogeneity gains are sampled from a uniform distribution in $a_{\text{min}},a_{\text{max}}]$ and applied with probability $0.8$ to each sample in the batch. If $|a|<0.05$, we set it to 0. Gains are clipped so that the maximum absolute value of the waveform does not exceed $1$. We anneal gain range from $(a_{\text{min}}=0,a_{\text{max}}=3)$ to no gain $(1,1)$ over training using a piecewise cosine schedule. Gain boundaries are defined as: $a_{\text{min}}(k)=1-\cdot C(k)$ and $a_{\text{max}}(k)=1+2\cdot C(k)$, where $C(k)$ is $1$ if $k\leq 0.2K$, 0 if $k\geq 0.9K$, and follows a cosine decay $C(k)=\frac{1}{2}(1+\cos(\pi\cdot\frac{k-0.2K}{0.7K}))$ for the intermediate steps.
4 Experiments and Results
Results are displayed in Tables 1 and 2. Best results are in bold, second best are underlined, and arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) is better. Audio examples and supplementary material are available online  [^2] https://bernardo-torres.github.io/projects/linear-cae.
Figure 3: Oracle Music Source Separation via latent arithmetic.
4.1 Experimental setup
Data: The training corpus is a large-scale music/speech dataset compiled from MTG-Jamendo [32], MoisesDB [33], M4Singer[34], DNS-Challenge [35], and E-GMD [36]. Tracks are sampled with weights $(60,20,9,8,3)$ respectively, and a segment of $2$ seconds is randomly cropped from each track. Segments are converted to mono and resampled to $44.1$ kHz.
Baselines: We report metrics using the publicly available weights of Music2Latent (M2L-Pub) [4] . As a fairer baseline, we retrain M2L-Pub on the same data as our model, including the random gains and artificial mixtures for data augmentation, but without our proposed implicit homogeneity and additivity conditioning strategies (M2L). We also report metrics for the public weights from Stable Audio $1.0$ VAE [11] (SA-VAE), serving as another autoencoder baseline with a different architecture and training procedure.
Table 1: Additivity and oracle Music Source Separation using latent space arithmetic on the MUSDB18-HQ dataset.
Additivity Reconstruction Separation
Decoder Encoder Mix Bass Drums Other Vocals
Model MSS $\downarrow$ SNR $\uparrow$ Error $\downarrow$ MSS $\downarrow$ SNR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$
M2L-Pub [4] $5.01$ 0.79-0.79 $2.82$ $1.16$ 1.29-1.29 $95.34$ 14.85-14.85 $28.08$ 16.30-16.30 $6.78$ 16.11-16.11 $5.10$ 18.16-18.16
SA-VAE [11] $5.38$ 12.58-12.58 $1.71$ $\mathbf{0.57}$ $\mathbf{4.03}$ $323.26$ $\underline{-2.81}$ $87.80$ $\underline{-3.11}$ $3.65$ $\underline{-4.10}$ $3.08$ $\underline{-4.48}$
M2L $5.21$ 0.48-0.48 $2.73$ $0.97$ $0.10$ $39.56$ 11.11-11.11 $12.26$ 11.95-11.95 $5.30$ 11.81-11.81 $3.49$ 12.56-12.56
Lin-CAE $\mathbf{0.99}$ $\mathbf{1.22}$ $\mathbf{0.60}$ $1.00$ $0.21$ $\mathbf{1.58}$ $\mathbf{-0.59}$ $\mathbf{1.16}$ $\mathbf{-0.28}$ $\mathbf{1.21}$ $\mathbf{-1.82}$ $\mathbf{1.23}$ $\mathbf{-1.18}$
Ablations
- Additivity $\underline{1.68}$ $\underline{0.30}$ $2.07$ $1.39$ 0.06-0.06 $\underline{6.61}$ 7.12-7.12 $\underline{2.37}$ 6.11-6.11 $3.44$ 8.52-8.52 $\underline{2.87}$ 12.18-12.18
- Homogeneity $4.24$ 0.89-0.89 $\underline{1.42}$ $\underline{0.96}$ $\underline{0.43}$ $23.61$ 5.96-5.96 $7.84$ 5.14-5.14 $\underline{2.98}$ 8.30-8.30 $3.08$ 7.95-7.95
Table 2: MusicCaps Reconstruction and Homogeneity (Hom.).
Reconstruction Dec-Hom. Enc-Hom.
Model MSS $\downarrow$ SNR $\uparrow$ MSS $\downarrow$ SNR $\uparrow$ Error $\downarrow$ KAD $\downarrow$
M2L-Pub [4] $1.14$ $1.85$ $2.52$ 4.69-4.69 $12.13$ $\mathbf{5.69}$
SA-VAE [11] $\mathbf{0.72}$ $\mathbf{7.32}$ $3.03$ $\underline{-1.27}$ $\underline{4.59}$ $6.27$
M2L $\underline{0.98}$ $3.09$ $\underline{2.27}$ 2.30-2.30 $8.52$ $6.53$
Lin-CAE $1.01$ $\underline{3.19}$ $\mathbf{1.37}$ $\mathbf{0.86}$ $\mathbf{0.69}$ $\underline{6.19}$
4.2 Reconstruction Quality
We evaluate the reconstruction quality of the AEs on 2-second audio chunks ($0.5$s overlap) from the MusicCaps dataset [37]. We reconstruct the full track using overlap-add and compute three metrics: (I) The Signal to Noise Ratio (SNR) on the waveform. (II) A Multi-Scale Spectral distance (MSS) between the original and reconstructed log-mel spectrograms using multiple resolutions3338080 mel bands and hop lengths of $\approx$ $10$, $25$, and $50$ ms.. (III) The Kernel Audio Distance (KAD) [38] between the embedding distributions of the original and reconstructed tracks [^4] Embedding distribution metrics have become a standard way of evaluating generative models, so we include it for completeness. KAD is computed using half of the reconstructed tracks, using the other half as the reference set. We use the kadtk library with the LAION-CLAP model..
Table 2 summarizes the results. Our retrained baseline (M2L) already improves on the public M2L-Pub weights [4], likely due to our distinct training data, batch size and augmentations. The linearized model (Lin-CAE) remains notably comparable on MSS and improves further on SNR, indicating that our method does not degrade reconstruction fidelity. Compared to SA-VAE, Lin-CAE is slightly worse, a similar finding to [4, 26]. We note that SA-VAE uses different data and a two-stage training procedure that optimizes both reconstruction and adversarial losses, while Lin-CAE is trained with a single denoising objective. KAD scores for Lin-CAE are better than our retrained M2L-Pub model and SA-VAE , but are worse than the public M2L weights (possibly reflecting the different training data distribution).
4.3 Homogeneity
We use the same data and setup as for reconstruction, but we additionally assess the model’s ability to preserve scaling by drawing a random scalar $\color[rgb]{1,0,1}a$}} for each track $x$ and measuring the effect on the output of both the encoder and decoder. We measure (I) Encoder Homogeneity (Enc-Hom.)as the relative $L_{2}$ error between the scaled input’s latent and the latent of the scaled input: $\|\operatorname{Enc}_{\theta}({{\color[rgb]{1,0,1}a}}\cdot x)-{{\color[rgb]{1,0,1}a}}\cdot\operatorname{Enc}_{\theta}(x)\|_{2}/\|{{\color[rgb]{1,0,1}a}}\cdot\operatorname{Enc}_{\theta}(x)\|_{2}$. We employ a relative error to normalize for the smaller latent norms produced by models trained with homogeneity, which would make a standard $L_{2}$ comparison misleading; (II) Decoder Homogeneity (Dec-Hom.)as the SNR and MSS between ${{\color[rgb]{1,0,1}a}}\cdot\operatorname{Dec}_{\theta}({\mathbf{z}}_{x})$ and $\operatorname{Dec}_{\theta}({{\color[rgb]{1,0,1}a}}\cdot{\mathbf{z}}_{x})$.
Results are shown in Table 2, where it is shown that Lin-CAE achieves much better decoder homogeneity (Table 2) properties compared to both baselines. For all baselines, we see a significant degradation in reconstruction metrics when testing the effect of random gains, while our model is much more robust. A high degradation in the log-domain MSS indicates that the scaling of the latent translates to significant timbral changes in the output. Lin-CAE’s MSS degrades less ($1.01\to 1.37$) than the baselines (M2L: $0.98\to 2.27$ and SA-VAE: $0.72\to 3.03$). The encoder also exhibits significantly lower homogeneity errors compared to all baselines.
4.4 Additivity
Additivity is evaluated on four source mixtures from the Musdb18-HQ test set [39], where a mixture is : $\text{mix}=\sum^{4}_{i=1}s_{i}$. $s_{i}$ correspond to either vocals, bass, drum or other. We measure Encoder Additivity Error as the deviation of a mixture’s latent from the sum of its source latents, calculated as the relative $L_{2}$ error: $\|\operatorname{Enc}_{\theta}(\text{mix})-\sum_{i}\operatorname{Enc}_{\theta}(s_{i})\|_{2}/\|\operatorname{Enc}_{\theta}(\text{mix})\|_{2}$. For Decoder Additivity (Composability), we test the ability to reconstruct a mixture from the sum of its source latents by SNR and MSS between the decoded sum of latents, $\operatorname{Dec}_{\theta}(\sum_{i}\operatorname{Enc}_{\theta}(s_{i}))$, to the autoencoded mixture, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix}))$.
Our results (Table 1) show that Lin-CAE achieves a very low MSS ($0.99$) compared to the baselines (>5>5), indicating that the summation in the latent space is very close to the autoencoded mixture, while the baselines produce a significantly degraded output. We encourage the reader to listen to the audio examples, as this is very clearly perceptible. The encoder also exhibits significantly lower additivity errors. Paired with the results from encoder homogeneity, this suggests that our implicit conditioning strategy encourages the entire AE mapping to become more linear, not just the decoder.
4.5 Music source separation via latent arithmetic
We perform source separation via latent arithmetic on the Musdb18-HQ test set. A source $\hat{s}_{i}$ is estimated by subtracting the latent representation of its accompaniment from the latent of the full mixture: $\hat{s}_{i}=\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix})-\operatorname{Enc}_{\theta}(\text{mix}-s_{i}))$. Figure 3 illustrates this process. The quality of the separated source is evaluated against the autoencoded ground-truth source, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(s_{i}))$, using Scale-invariant Signal to Distortion Ratio (SI-SDR) [40] and MSS.
Table 1 shows that Lin-CAE significantly outperforms the baselines in latent arithmetic/oracle source separation tasks, achieving higher SI-SDR and lower MSS scores by a large margin for every instrument. Notably, both separation metrics are of the same order of magnitude as the reconstruction values for the mix, which can be seen as an upper bound for the model. That is not the case, however, for any of the baselines, with the least significant gap for SA-VAE. We note that during training, we have only shown positive gains to the model, while during separation, we are effectively applying negative gains when subtracting latents. This further indicates the robustness of the learned linear structure.
Ablations: We retrain Lin-CAE with only one of the two properties (homogeneity or additivity) and report source separation results in Table 1. Performance in every metric degrades significantly when only one of the two properties is applied. Surprisingly, training with only homogeneity seems to help the additivity and source separation much more as a byproduct than training with only additivity.
5 Conclusion
We introduce a simple method to induce an approximately linear latent space in audio autoencoders during training. Using implicit conditioning via data augmentation, our approach enforces the properties of a linear map (homogeneity and additivity) without necessitating any modifications to the architecture or loss function. Experiments with Consistency Autoencoders (CAEs) show that linearization preserves high-quality reconstruction and can be used for source separation via latent arithmetic.
Our approach enables more interpretable and controllable audio generation. Future work could extend this method to improved CAE architectures [26] and downstream tasks like generative source separation [41]. Direct, low-level manipulations in compressed space promise new possibilities for audio editing and efficient signal processing.

### v2

@@
 Modern Autoencoders can achieve excellent reconstruction quality at high compression rates at the expense of complex, entangled latent spaces.
-For applications where input space manipulation is desirable from within the compressed space, recent research has proposed either task-specific adaptors for pre-trained models or redesigning the autoencoder to preserve key structural properties, such as equivariance to spatial transformations [1, 2, 3]).
+For applications where input space manipulation is desirable from within the compressed space, recent research has proposed either task-specific adaptors for pre-trained models or redesigning the autoencoder to preserve key structural properties, such as equivariance to spatial transformations [kouzelisEQVAEEquivarianceRegularized2025, skorokhodovImprovingDiffusabilityAutoencoders2025, zhouAliasfreeLatentDiffusion2025].
 This work follows the latter approach.
@@
 The method employs implicit regularization through data augmentation, without modifying the model architecture or objective.
-The approach is demonstrated with the Music2Latent architecture [4], a Consistency Autoencoder (CAE) that achieves high-quality, single-step reconstruction and a $64\times$ compression rate for $44.1$ kHz audio.
+The approach is demonstrated with the Music2Latent architecture [pasiniMusic2LatentConsistencyAutoencoders2024], a Consistency Autoencoder (CAE) that achieves high-quality, single-step reconstruction and a $64\times$ compression rate for $44.1$ kHz audio.
 Figure 1: In a linear decoder, applying a gain to the latent vector scales the output by the same gain (homogeneity), and summing latents corresponds to a sum in the audio domain (additivity).
@@
 2.1 Audio processing in the latent space
-There is a growing interest in manipulating audio by training task-specific modules which operate directly in the latent space of a pretrained AE, with applications in source separation [5], speech enhancement [6, 7], upsampling and upmixing [8], filtering [9], and generative modelling [10, 11, 12].
-Some works retrain AEs for specific objectives, such as source disentanglement [13] or separation via latent masking [14].
+There is a growing interest in manipulating audio by training task-specific modules which operate directly in the latent space of a pretrained AE, with applications in source separation [mancusiUnsupervisedSourceSeparation2021, postolacheLatentAutoregressiveSource2023, bindiUnsupervisedComposableRepresentations2024], speech enhancement [omranDisentanglingSpeechSurroundings2023, liSpeechEnhancementUsing2025], upsampling and upmixing [braliosLearningUpsampleUpmix2025], filtering [braliosRebottleneckLatentRestructuring2025], and generative modelling [schneiderMousaiEfficientTexttomusic2024, evansStableAudioOpen2025, nistalDiffariffMusicalAccompaniment2024].
+Some works retrain AEs for specific objectives, such as source disentanglement [bieLearningSourceDisentanglement2025] or separation via latent masking [tzinisTwostepSoundSource2020].
 In contrast, our work focuses on training an AE to have desirable, task-agnostic structural properties.
@@
 2.2 Diffusion and Consistency models
-Denoising Diffusion Probabilistic Models [15] and score-based models [16] have achieved great success in generative modeling, where the goal is to estimate the underlying data distribution from samples.
+Denoising Diffusion Probabilistic Models [hoDenoisingDiffusionProbabilistic2020a] and score-based models [songScorebasedGenerativeModeling2021] have achieved great success in generative modeling, where the goal is to estimate the underlying data distribution from samples.
 These models define a forward process that gradually adds noise to data and learn a reverse process to denoise it using a neural network.
 Sampling typically requires an iterative procedure with many steps to generate a clean sample.
-Consistency Models [17] accelerate this process by mapping any point along the trajectory defined by the probability flow ordinary differential equation [16] directly to the origin (the clean data point), enabling single-step generation.
+Consistency Models [songConsistencyModels2023] accelerate this process by mapping any point along the trajectory defined by the probability flow ordinary differential equation [songScorebasedGenerativeModeling2021] directly to the origin (the clean data point), enabling single-step generation.
 When trained from scratch, the process is called Consistency Training (CT), in which a ”student” denoiser network ($f_{\theta}$) is trained to match the output of a ”teacher” ($f_{\theta^{-}}$), which itself denoises a less corrupted version of the same data point.
-In improved Consistency Training (iCT) [18], the teacher is updated with the same parameters $\theta$ as the student but detached from the computational graph ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$).
+In improved Consistency Training (iCT) [songImprovedTechniquesTraining2023], the teacher is updated with the same parameters $\theta$ as the student but detached from the computational graph ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$).
 2.3 High-Fidelity Audio Autoencoders
-In audio, popular AEs include Neural Audio Codecs [19], which compress audio into a set of discrete tokens, and Variational Autoencoders [20], which have seen significant success in generative modeling.
-However, both frameworks require complex, multi-stage training with adversarial objectives [21] to achieve high-quality reconstructions.
-Diffusion Autoencoders [22, 23, 24, 25, 10] enable high-fidelity reconstruction by replacing the deterministic decoder with a conditional diffusion model.
-Notable examples in audio include [10] and Music2Latent [4, 26].
-A specific instance of DiffAEs is the CAE[4, 26], where the decoder is a CM, enabling decoding in one step.
-DiffAEs can be trained with a single diffusion/CT objective [22, 24, 4].
-High-quality decoding has been linked to the capacity of sampling details at inference-time, since the decoder acts as a denoiser instead of an upsampler, which also reduces the amount of unnecessary information to be encoded in the latent [23, 25].
+In audio, popular AEs include Neural Audio Codecs [defossezHighFidelityNeural2023], which compress audio into a set of discrete tokens, and Variational Autoencoders [kingmaAutoencodingVariationalBayes2014], which have seen significant success in generative modeling.
+However, both frameworks require complex, multi-stage training with adversarial objectives [esserTamingTransformersHighresolution2021] to achieve high-quality reconstructions.
+Diffusion Autoencoders [preechakulDiffusionAutoencodersMeaningful2022, birodkarSampleWhatYou2024, chenDiffusionAutoencodersAre2025, zhaoEpsilonvaeDenoisingVisual2024, schneiderMousaiEfficientTexttomusic2024] enable high-fidelity reconstruction by replacing the deterministic decoder with a conditional diffusion model.
+Notable examples in audio include [schneiderMousaiEfficientTexttomusic2024] and Music2Latent [pasiniMusic2LatentConsistencyAutoencoders2024, pasiniMusic2Latent2AudioCompression2025].
+A specific instance of DiffAEs is the CAE[pasiniMusic2LatentConsistencyAutoencoders2024, pasiniMusic2Latent2AudioCompression2025], where the decoder is a CM, enabling decoding in one step.
+DiffAEs can be trained with a single diffusion/CT objective [preechakulDiffusionAutoencodersMeaningful2022, chenDiffusionAutoencodersAre2025, pasiniMusic2LatentConsistencyAutoencoders2024].
+High-quality decoding has been linked to the capacity of sampling details at inference-time, since the decoder acts as a denoiser instead of an upsampler, which also reduces the amount of unnecessary information to be encoded in the latent [birodkarSampleWhatYou2024, zhaoEpsilonvaeDenoisingVisual2024].
 3 Method
@@
 $\operatorname{Dec}_{\theta}({\mathbf{z}}_{u}+{\mathbf{z}}_{v})\approx\operatorname{Dec}_{\theta}({\mathbf{z}}_{u})+\operatorname{Dec}_{\theta}({\mathbf{z}}_{v}),\quad\forall{\mathbf{z}}_{u},{\mathbf{z}}_{v}\in\mathcal{Z}$ (2)
-We postulate that approximate linearity can be achieved by a simple data augmentation scheme, without the need to change model or loss function.
-Our approach is in theory model-agnostic and can be applied to any AE, and in this work we apply it on a CAE architecture for audio [4], which offers $64\times$ compression and high-quality reconstruction with a single training objective.
+We postulate that approximate linearity can be achieved through a simple data augmentation scheme, without the need to modify the model or loss function.
+The intuition is that by performing latent space operations (scaling and addition) and asking the decoder to recover the corresponding audio signal, it learns to behave linearly.
+Our approach is, in theory, model-agnostic and can be applied to any AE.
+In this work, we apply it to a CAE architecture for audio [pasiniMusic2LatentConsistencyAutoencoders2024], which offers $64\times$ compression and high-quality reconstruction with a single training objective.
 3.1 Model architecture
-Our model (Lin-CAE) follows the Music2Latent architecture [4], illustrated in Figure 2(2(a)). We recall the main components here for clarity, but we refer the reader to [4] for the full details.
-CAE Denoiser representation space: We use a complex-valued Short-Time Fourier Transform (STFT) followed by the invertible amplitude scaling ${\operatorname{Amp}}:{\mathbb{C}}\to{\mathbb{C}}=\beta|c|^{\alpha}e^{i\angle(c)}$ (with $\alpha=0.65$ and $\beta=0.34$) from previous work in complex STFT diffusion [27, 28, 4], which scales the amplitudes in the STFT to roughly [1,1][-1,1] while boosting the high frequencies.
+Our model (Lin-CAE) follows the Music2Latent architecture [pasiniMusic2LatentConsistencyAutoencoders2024], illustrated in Figure 2(2(a)).
+We recall the main components here for clarity, but we refer the reader to [pasiniMusic2LatentConsistencyAutoencoders2024] for the full details.
+CAE Denoiser representation space: We use a complex-valued Short-Time Fourier Transform (STFT) followed by the invertible amplitude scaling ${\operatorname{Amp}}:{\mathbb{C}}\to{\mathbb{C}}=\beta|c|^{\alpha}e^{i\angle(c)}$ (with $\alpha=0.65$ and $\beta=0.34$) from previous work in complex STFT diffusion [richterSpeechEnhancementDereverberation2023, zhuEdmsoundSpectrogramBased2023, pasiniMusic2LatentConsistencyAutoencoders2024], which scales the amplitudes in the STFT to roughly [1,1][-1,1] while boosting the high frequencies.
 We define ${\operatorname{T}}(x)={\operatorname{Amp}}(\operatorname{STFT}(x))$ as the transform which maps from waveform to CM input space.
 Decoder: $\operatorname{Dec}_{\theta}$ is composed of a U-Net-based denoiser $f_{\theta}$} and the inverse transform ${\operatorname{T}}^{-1}({\mathbf{x}})$.
-The denoiser reconstructs ${\mathbf{x}}={\operatorname{T}}(x)$ from a noise-corrupted version ${\mathbf{x}}_{\sigma}={\mathbf{x}}+\sigma\bm{\epsilon}$ ($\bm{\epsilon}\sim\mathcal{N}(0,I)$), conditioned on the latent ${\mathbf{z}}_{x}$ and the noise level $\sigma$.
+The denoiser reconstructs ${\mathbf{x}}={\operatorname{T}}(x)$ from a noise-corrupted version ${\mathbf{x}}_{\sigma}={\mathbf{x}}+\sigma\cdot\bm{\epsilon}$ ($\bm{\epsilon}\sim\mathcal{N}(0,I)$), conditioned on the latent ${\mathbf{z}}_{x}$ and the noise level $\sigma$.
 ${\mathbf{z}}_{x}$ is upsampled with a dedicated network mirroring the U-Net’s upsampling block and added to each layer of the U-Net.
 Noise level information is added to every layer via positional embeddings.
-$f_{\theta}$} is parameterized by a noise prediction network $F_{\theta}$} with a skip connection weighted by coefficients ($c_{\text{skip}},c_{\text{out}}$) that depend on $\sigma$ [29], which additionally enforces the boundary condition necessary for CT [17, 18, 4]:
+$f_{\theta}$} is parameterized by a noise prediction network $F_{\theta}$} with a skip connection weighted by coefficients ($c_{\text{skip}},c_{\text{out}}$) that depend on $\sigma$ [karrasElucidatingDesignSpace2022], which additionally enforces the boundary condition necessary for CT [songConsistencyModels2023, songImprovedTechniquesTraining2023, pasiniMusic2LatentConsistencyAutoencoders2024]:
 ${f_{\theta}}({\mathbf{x}}_{\sigma},\sigma,{\mathbf{z}}_{x})=c_{\text{skip}}(\sigma){\mathbf{x}}_{\sigma}+c_{\text{out}}(\sigma){F_{\theta}}({\mathbf{x}}_{\sigma},\sigma,{\mathbf{z}}_{x})$ (3)
@@
 3.2 Learning Linearity in CAEs via Implicit Regularization
-Implicit homogeneity: Inspired by [1], we apply a random positive gain ($\color[rgb]{1,0,1}a$}}) to the latent and task the decoder to reconstruct a scaled version of the input (${{\color[rgb]{1,0,1}a}}x$).
+Implicit homogeneity: Inspired by [kouzelisEQVAEEquivarianceRegularized2025], we apply a random positive gain ($\color[rgb]{1,0,1}a$}}) to the latent and task the decoder to denoise a scaled version of the input (${{\color[rgb]{1,0,1}a}}\cdot x$).
 The adapted CT training, depicted in Figure 2(2(b)), uses two parallel pathways.
 $\operatorname{Enc}_{\theta}$ encodes unscaled $x$ to obtain ${\mathbf{z}}_{x}$.
-Then, the denoiser ($f_{\theta}$}) receives a noisy scaled input (${\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma\bm{\epsilon}$) conditioned on the scaled latent (${{\color[rgb]{1,0,1}a}}{\mathbf{z}}_{x}$).
+Then, the denoiser ($f_{\theta}$}) receives a noisy scaled input (${\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot x)+\sigma\cdot\bm{\epsilon}$) conditioned on the scaled latent (${{\color[rgb]{1,0,1}a}}\cdot{\mathbf{z}}_{x}$).
 $\color[rgb]{1,0,1}a$}} is never provided as an explicit input to any part of the model, so $\operatorname{Dec}_{\theta}$ must learn to infer the correct output scale from the magnitude of the conditioning latent.
 If ${{\color[rgb]{1,0,1}a}}=1$, we recover the original CAE training.
-Implicit additivity: Inspired by Mixit [30], we augment our data by creating artificial mixes from pairs of elements $(u,v)$ randomly selected from the training set.
-When decoding, we replace the true latent of the artificial mix ($\operatorname{Enc}_{\theta}(u+v)$) with the sum of the latents of the individual signals: ${\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}}\leftarrow({\color[rgb]{0,0,1}{\mathbf{z}}_{u}+{\mathbf{z}}_{v}})$, where $({\mathbf{z}}_{u}=\operatorname{Enc}_{\theta}(u),{\mathbf{z}}_{v}=\operatorname{Enc}_{\theta}(v))$, as illustrated in Figure 2(2(c)).
-The denoiser is thus conditioned on $\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}$} and tasked with denoising ${\operatorname{T}}({{\color[rgb]{1,0,1}a}}(u+v))+\sigma\bm{\epsilon})$.
-We create mixtures from each training batch by concatenating it with a version of itself that has been circularly shifted by one position and summed with the original (Figure 2(2(d))).
+Implicit additivity: Inspired by Mixit [wisdomUnsupervisedSoundSeparation2020], we augment our data by creating artificial mixes from pairs of elements $(u,v)$ randomly selected from the training set.
+When decoding, we replace the true latent of the artificial mix ($\operatorname{Enc}_{\theta}(u+v)$) with the average of the latents of the individual signals: ${\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}}\leftarrow 0.5\cdot({\mathbf{z}}_{u}+{\mathbf{z}}_{v})$, where $({\mathbf{z}}_{u}=\operatorname{Enc}_{\theta}(u),{\mathbf{z}}_{v}=\operatorname{Enc}_{\theta}(v))$, as illustrated in Figure 2(2(c)).
+The denoiser is thus conditioned on $\color[rgb]{0,0,1}{\mathbf{z}}^{\prime}$} and tasked with denoising ${\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot 0.5\cdot(u+v))+\sigma\cdot\bm{\epsilon})$.
+We augment the training batch with mixtures by concatenating it with a version of itself that has been circularly shifted by one position and averaged with the original (Figure 2(2(d))).
 3.3 Training
-With the modifications described in Section 3.2 highlighted in color, we train the model under the CT objective $\mathcal{L}_{\text{CT}}$ [17, 18, 4], which minimizes the distance between the denoised output of a teacher ($f_{\theta^{-}}$) and student ($f_{\theta}$) models ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$) for noisy inputs at two different noise levels, $\sigma_{t_{1}}$ and $\sigma_{t_{2}}$.
-$\begin{split}\mathcal{L}_{\text{CT}}=\mathbb{E}_{{\mathbf{x}},\bm{\epsilon},t_{1},{{\color[rgb]{1,0,1}a}}}\bigg[\lambda\cdot d\Big(f_{\theta}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma_{t_{2}}\bm{\epsilon},\sigma_{t_{2}},{{\color[rgb]{1,0,1}a}}{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime}),\\ f_{\theta^{-}}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}x)+\sigma_{t_{1}}\bm{\epsilon},\sigma_{t_{1}},{{\color[rgb]{1,0,1}a}}{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime})\Big)\bigg],\end{split}$ (4)
-where $\color[rgb]{0,0,1}{\mathbf{z}}$$^{\prime}$ is either ${\mathbf{z}}_{x}$ or ${\mathbf{z}}_{u}+{\mathbf{z}}_{v}$ depending on whether the input waveform is sampled from the training set or is an artificial mixture ($x=u+v$), $\bm{\epsilon}\sim\mathcal{N}(0,I)$ is a shared noise direction, $\lambda(\sigma_{t_{1}},\sigma_{t_{2}})=\frac{1}{\sigma_{t_{2}}-\sigma_{t_{1}}}$ is a weighting function designed to give higher weights to smaller noise increments [18].
-We sample time steps $t_{1}\in[0,1]$ and corresponding noise levels $\sigma_{t_{1},t_{2}}$ following [4], with $t_{2}=t_{1}+\Delta_{t_{k}}$, where $\Delta_{t_{k}}$ depends on the training step $k$ and decays over training ($\Delta t_{k}=\Delta t_{0}^{\frac{k}{K}(e_{K}-1)+1}$, $\Delta t_{0}=0.1$, $e_{K}=2.0$).
-$d$ is the Pseudo-Huber loss [31]: $d(x,y)=\sqrt{|x-y|^{2}+c^{2}}-c$, with c=5.4e4c=5.4e-4.
+With the modifications described in Section 3.2 highlighted in color, we train the model under the CT objective $\mathcal{L}_{\text{CT}}$ [songConsistencyModels2023, songImprovedTechniquesTraining2023, pasiniMusic2LatentConsistencyAutoencoders2024], which minimizes the distance between the denoised output of a teacher ($f_{\theta^{-}}$) and student ($f_{\theta}$) models ($\theta^{-}\leftarrow\text{stopgrad}(\theta)$) for noisy inputs at two different noise levels, $\sigma_{t_{1}}$ and $\sigma_{t_{2}}$.
+$\begin{split}\mathcal{L}_{\text{CT}}=\mathbb{E}_{{\mathbf{x}},\bm{\epsilon},t_{2},{{\color[rgb]{1,0,1}a}}}\bigg[\lambda\cdot d\Big(f_{\theta}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot x)+\sigma_{t_{2}}\cdot\bm{\epsilon},\sigma_{t_{2}},{{\color[rgb]{1,0,1}a}}\cdot{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime}),\\ f_{\theta^{-}}&({\operatorname{T}}({{\color[rgb]{1,0,1}a}}\cdot x)+\sigma_{t_{1}}\cdot\bm{\epsilon},\sigma_{t_{1}},{{\color[rgb]{1,0,1}a}}\cdot{{\color[rgb]{0,0,1}{\mathbf{z}}}}^{\prime})\Big)\bigg],\end{split}$ (4)
+where $\color[rgb]{0,0,1}{\mathbf{z}}$$^{\prime}$ is either ${\mathbf{z}}_{x}$ or $0.5\cdot({\mathbf{z}}_{u}+{\mathbf{z}}_{v})$ depending on whether the input waveform is sampled from the training set or is an artificial mixture ($x=0.5\cdot(u+v$)), $\bm{\epsilon}\sim\mathcal{N}(0,I)$ is a shared noise direction, $\lambda(\sigma_{t_{1}},\sigma_{t_{2}})=\frac{1}{\sigma_{t_{2}}-\sigma_{t_{1}}}$ is a weighting function designed to give higher weights to smaller noise increments [songImprovedTechniquesTraining2023].
+We sample time steps $t_{2}\in[0,1]$ and corresponding noise levels $\sigma_{t_{1},t_{2}}$ following [pasiniMusic2LatentConsistencyAutoencoders2024], with $t_{1}=\max(t_{2}-\Delta_{t_{k}},0)$, where $\Delta_{t_{k}}$ depends on the training step $k$ and decays over training ($\Delta t_{k}=\Delta t_{0}^{\frac{k}{K}(e_{K}-1)+1}$, $\Delta t_{0}=0.1$, $e_{K}=2.0$).
+$d$ is the Pseudo-Huber loss [charbonnierDeterministicEdgepreservingRegularization1997]: $d(x,y)=\sqrt{|x-y|^{2}+c^{2}}-c$, with c=5.4e4c=5.4e-4.
 Training details: We train all models for 800800K steps with a batch size of $20$ before mixture creation (final batch size of $40$).
@@
 4.1 Experimental setup
-Data: The training corpus is a large-scale music/speech dataset compiled from MTG-Jamendo [32], MoisesDB [33], M4Singer[34], DNS-Challenge [35], and E-GMD [36].
+Data: The training corpus is a large-scale music/speech dataset compiled from MTG-Jamendo [bogdanovMTGjamendoDatasetAutomatic2019], MoisesDB [pereiraMoisesDBDatasetSource2023], M4Singer[zhangM4SingerMultistyleMultisinger2022], DNS-Challenge [dubeyICASSP2023Deep2023], and E-GMD [callenderImprovingPerceptualQuality2020].
 Tracks are sampled with weights $(60,20,9,8,3)$ respectively, and a segment of $2$ seconds is randomly cropped from each track.
-Segments are converted to mono and resampled to $44.1$ kHz.
-Baselines: We report metrics using the publicly available weights of Music2Latent (M2L-Pub) [4] .
+Segments are converted to mono by averaging stereo channels and resampled to $44.1$ kHz.
+Baselines: We report metrics using the publicly available weights of Music2Latent (M2L-Pub) [pasiniMusic2LatentConsistencyAutoencoders2024] .
 As a fairer baseline, we retrain M2L-Pub on the same data as our model, including the random gains and artificial mixtures for data augmentation, but without our proposed implicit homogeneity and additivity conditioning strategies (M2L).
-We also report metrics for the public weights from Stable Audio $1.0$ VAE [11] (SA-VAE), serving as another autoencoder baseline with a different architecture and training procedure.
+We also report metrics for the public weights from Stable Audio $1.0$ VAE [evansStableAudioOpen2025] (SA-VAE), serving as another autoencoder baseline with a different architecture and training procedure.
 Table 1: Additivity and oracle Music Source Separation using latent space arithmetic on the MUSDB18-HQ dataset.
@@
 Model MSS $\downarrow$ SNR $\uparrow$ Error $\downarrow$ MSS $\downarrow$ SNR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$ MSS $\downarrow$ SI-SDR $\uparrow$
-M2L-Pub [4] $5.01$ 0.79-0.79 $2.82$ $1.16$ 1.29-1.29 $95.34$ 14.85-14.85 $28.08$ 16.30-16.30 $6.78$ 16.11-16.11 $5.10$ 18.16-18.16
-SA-VAE [11] $5.38$ 12.58-12.58 $1.71$ $\mathbf{0.57}$ $\mathbf{4.03}$ $323.26$ $\underline{-2.81}$ $87.80$ $\underline{-3.11}$ $3.65$ $\underline{-4.10}$ $3.08$ $\underline{-4.48}$
+M2L-Pub [pasiniMusic2LatentConsistencyAutoencoders2024] $5.01$ 0.79-0.79 $2.82$ $1.16$ 1.29-1.29 $95.34$ 14.85-14.85 $28.08$ 16.30-16.30 $6.78$ 16.11-16.11 $5.10$ 18.16-18.16
+SA-VAE [evansStableAudioOpen2025] $5.38$ 12.58-12.58 $1.71$ $\mathbf{0.57}$ $\mathbf{4.03}$ $323.26$ $\underline{-2.81}$ $87.80$ $\underline{-3.11}$ $3.65$ $\underline{-4.10}$ $3.08$ $\underline{-4.48}$
 M2L $5.21$ 0.48-0.48 $2.73$ $0.97$ $0.10$ $39.56$ 11.11-11.11 $12.26$ 11.95-11.95 $5.30$ 11.81-11.81 $3.49$ 12.56-12.56
@@
 Model MSS $\downarrow$ SNR $\uparrow$ MSS $\downarrow$ SNR $\uparrow$ Error $\downarrow$ KAD $\downarrow$
-M2L-Pub [4] $1.14$ $1.85$ $2.52$ 4.69-4.69 $12.13$ $\mathbf{5.69}$
-SA-VAE [11] $\mathbf{0.72}$ $\mathbf{7.32}$ $3.03$ $\underline{-1.27}$ $\underline{4.59}$ $6.27$
+M2L-Pub [pasiniMusic2LatentConsistencyAutoencoders2024] $1.14$ $1.85$ $2.52$ 4.69-4.69 $12.13$ $\mathbf{5.69}$
+SA-VAE [evansStableAudioOpen2025] $\mathbf{0.72}$ $\mathbf{7.32}$ $3.03$ $\underline{-1.27}$ $\underline{4.59}$ $6.27$
 M2L $\underline{0.98}$ $3.09$ $\underline{2.27}$ 2.30-2.30 $8.52$ $6.53$
@@
 4.2 Reconstruction Quality
-We evaluate the reconstruction quality of the AEs on 2-second audio chunks ($0.5$s overlap) from the MusicCaps dataset [37].
+We evaluate the reconstruction quality of the AEs on 2-second audio chunks ($0.5$s overlap) from the MusicCaps dataset [agostinelliMusicLMGeneratingMusic2023].
 We reconstruct the full track using overlap-add and compute three metrics: (I) The Signal to Noise Ratio (SNR) on the waveform.
-(II) A Multi-Scale Spectral distance (MSS) between the original and reconstructed log-mel spectrograms using multiple resolutions3338080 mel bands and hop lengths of $\approx$ $10$, $25$, and $50$ ms..
-(III) The Kernel Audio Distance (KAD) [38] between the embedding distributions of the original and reconstructed tracks [^4] Embedding distribution metrics have become a standard way of evaluating generative models, so we include it for completeness.
+(II) A Multi-Scale Spectral distance (MSS) between the original and reconstructed log-mel spectrograms [^3] We use mel-spectrogram distances computed over multiple resolutions ($80$ mel bands and hop lengths of $\approx$ $10$, $25$, and $50$ ms) as a proxy for a phase-insensitive perceptual metric.
+.
+(III) The Kernel Audio Distance (KAD) [chungKADNoMore2025] between the embedding distributions of the original and reconstructed tracks [^4] Embedding distribution metrics have become a standard way of evaluating generative models, so we include it for completeness.
 KAD is computed using half of the reconstructed tracks, using the other half as the reference set.
@@
 Table 2 summarizes the results.
-Our retrained baseline (M2L) already improves on the public M2L-Pub weights [4], likely due to our distinct training data, batch size and augmentations.
+Our retrained baseline (M2L) already improves on the public M2L-Pub weights [pasiniMusic2LatentConsistencyAutoencoders2024], likely due to our distinct training data, batch size and augmentations.
 The linearized model (Lin-CAE) remains notably comparable on MSS and improves further on SNR, indicating that our method does not degrade reconstruction fidelity.
-Compared to SA-VAE, Lin-CAE is slightly worse, a similar finding to [4, 26].
+Compared to SA-VAE, Lin-CAE is slightly worse, a similar finding to [pasiniMusic2LatentConsistencyAutoencoders2024, pasiniMusic2Latent2AudioCompression2025].
 We note that SA-VAE uses different data and a two-stage training procedure that optimizes both reconstruction and adversarial losses, while Lin-CAE is trained with a single denoising objective.
@@
 4.4 Additivity
-Additivity is evaluated on four source mixtures from the Musdb18-HQ test set [39], where a mixture is : $\text{mix}=\sum^{4}_{i=1}s_{i}$.
+Additivity is evaluated on four source mixtures from the Musdb18-HQ test set [rafiiMUSDB18HQUncompressedVersion2019], where a mixture is : $\text{mix}=\sum^{4}_{i=1}s_{i}$.
 $s_{i}$ correspond to either vocals, bass, drum or other.
 We measure Encoder Additivity Error as the deviation of a mixture’s latent from the sum of its source latents, calculated as the relative $L_{2}$ error: $\|\operatorname{Enc}_{\theta}(\text{mix})-\sum_{i}\operatorname{Enc}_{\theta}(s_{i})\|_{2}/\|\operatorname{Enc}_{\theta}(\text{mix})\|_{2}$.
-For Decoder Additivity (Composability), we test the ability to reconstruct a mixture from the sum of its source latents by SNR and MSS between the decoded sum of latents, $\operatorname{Dec}_{\theta}(\sum_{i}\operatorname{Enc}_{\theta}(s_{i}))$, to the autoencoded mixture, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix}))$.
+For Decoder Additivity, we test the ability to reconstruct a mixture from the sum of its source latents by SNR/MSS between the decoded sum of latents, $\operatorname{Dec}_{\theta}(\sum_{i}\operatorname{Enc}_{\theta}(s_{i}))$, and the autoencoded mixture, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(\text{mix}))$.
 Our results (Table 1) show that Lin-CAE achieves a very low MSS ($0.99$) compared to the baselines (>5>5), indicating that the summation in the latent space is very close to the autoencoded mixture, while the baselines produce a significantly degraded output.
@@
 Figure 3 illustrates this process.
-The quality of the separated source is evaluated against the autoencoded ground-truth source, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(s_{i}))$, using Scale-invariant Signal to Distortion Ratio (SI-SDR) [40] and MSS.
+The quality of the separated source is evaluated against the autoencoded ground-truth source, $\operatorname{Dec}_{\theta}(\operatorname{Enc}_{\theta}(s_{i}))$, using Scale-invariant Signal to Distortion Ratio (SI-SDR) [lerouxSDRHalfbakedWell2019] and MSS.
 Table 1 shows that Lin-CAE significantly outperforms the baselines in latent arithmetic/oracle source separation tasks, achieving higher SI-SDR and lower MSS scores by a large margin for every instrument.
-Notably, both separation metrics are of the same order of magnitude as the reconstruction values for the mix, which can be seen as an upper bound for the model.
+The model was not trained to recover the exact phase, which is why the relatively low SDR values are observed for both reconstruction and separation.
+Notably, however, both separation metrics are of the same order of magnitude as the reconstruction values for the mix, which can be seen as an upper bound for the model.
 That is not the case, however, for any of the baselines, with the least significant gap for SA-VAE.
@@
 We introduce a simple method to induce an approximately linear latent space in audio autoencoders during training.
-Using implicit conditioning via data augmentation, our approach enforces the properties of a linear map (homogeneity and additivity) without necessitating any modifications to the architecture or loss function.
-Experiments with Consistency Autoencoders (CAEs) show that linearization preserves high-quality reconstruction and can be used for source separation via latent arithmetic.
+Using implicit conditioning via data augmentation, our approach enforces homogeneity and additivity (properties of a linear map) without modifying the architecture or objective.
+Experiments with Consistency Autoencoders show that linearization preserves high-quality reconstruction and can be used for source separation via latent arithmetic.
 Our approach enables more interpretable and controllable audio generation.
-Future work could extend this method to improved CAE architectures [26] and downstream tasks like generative source separation [41].
+Future work could extend this method to improved CAE architectures [pasiniMusic2Latent2AudioCompression2025] and downstream tasks such as generative source separation [marianiMultisourceDiffusionModels2024].
 Direct, low-level manipulations in compressed space promise new possibilities for audio editing and efficient signal processing.